# Natural Counterfactuals With Necessary Backtracking

### 02 Feb 2024 v1

@@ 2 Related Work
Counterfactual Explanations. A prominent approach in explainable AI is counterfactual explanation (Wachter et al., 2018; Dhurandhar et al., 2018; Mothilal et al., 2020; Barocas et al., 2020; Pawlowski et al., 2020a; Verma et al., 2020; Schut et al., 2021), on which our work is likely to have interesting bearings. Most works on this topic define some sense of minimal changes of an input sample with a predicted class $c$ such that adding the minimal changes into the input would make it be classified into another (more desirable) class. Although this paper does not discuss counterfactual explanations, our framework may well be used to define a novel notion of counterfactual explanation by requiring the counterfactual instances to be “natural” in our sense.
@@ 4.3 Distance Measure for Necessary Backtracking
4.4 Identifiablity of Natural Counterfactuals
@@ 6.2 MorphoMNIST
MorphoMNIST involves three variables $(t,i,x)$. As depicted in Fig. 3 (a), $t$ (digit stroke thickness) causes both $i$ (stroke intensity) and $x$ (images), with $i$ being the direct cause of $x$.
@@
In our experiments, mirroring those in Section 6.1, we incorporate two key changes. First, we utilize two advanced deep learning models, V-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023), for doing counterfactuals. Second, due to the absence of ground-truth SCM for assessing outcome error, we adopt the counterfactual effectiveness metric from (Ribeiro et al., 2023), as developed in (Monteiro et al., 2023). This involves training a predictor on the dataset to estimate parent values $(\hat{t},\hat{i})$ from a counterfactual image $x$ generated by learned model $p(x|t,i)$ with the input $(t,i)$, and then computing the absolute error $|t-\hat{t}|$ or $|i-\hat{i}|$ between the input $(t,i)$ and their predicted counterparts $(\hat{t},\hat{i})$ as inferred from counterfactual images.
@@ 7 Conclusion
To address the impracticality of hard interventions, we propose “natural counterfactuals,” a new counterfactual inference approach more applicable to real-world scenarios. This method integrates a naturalness constraint and necessary backtracking, forming an optimization framework that identifies feasible interventions, leading to outcomes better aligned with data distribution. We demonstrate the effectiveness of our approach in deep learning case studies. Future research will further investigate the use of natural counterfactuals in vital areas like healthcare, economics, and law.
@@ A.2 MorphoMNIST
The MorphoMNIST comes from (Pawlowski et al., 2020b), where there are 60000 images as training set and 10,000 images as test dataset. Fig. 8 (a) shows the causal graph for generating Morpho-MNIST; specifically, stroke thickness $t$ causes the brightness intensity $i$, and both thickness $t$ and intensity $i$ cause the digit $x$. Fig. 8 (b) show some samples from Morpho-MNIST. The ground-truth SCMs to generate the Morpho-MNIST are as follows:
@@ Appendix C Model Training
Our study focuses on counterfactual inference and we directly use two state-of-the-art deep-learning SCM models to learn SCMs among variables using a dataset, i.e., D-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023). Specifically, we use code of (Ribeiro et al., 2023) containing the implementation of D-SCM and H-SCM. Take Morpho-MNIST as an example, in both two models, normalizing flows are firstly trained to learn causal mechanisms for all variables except image $x$, i.e., $(t,i)$, and a conditional VAE is used to model image $x$ given its parents $(t,i)$. For D-SCM, the conditional VAE uses normal VAE framework, while H-SCM uses hierarchical VAE structure (Maaløe et al., 2019) to better capture the distribution of images.
@@ H.2.1 MorphoMNIST
We follow the experimental settings of simulation experiments in Sec. 6.1, except for two differences. One is that we use two state-of-the-art deep learning models, namely V-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023), as the backbones to learn counterfactuals. They use normalizing flows to learn causal relationships among $x$’s parent nodes, e.g., $(t,i)$ in MorphoMNIST. Further, to learn $p(x|t,i)$, notice that V-SCM uses VAE (Kingma & Welling, 2014) and HVAE (Maaløe et al., 2019). Another difference is that, instead of estimating the outcome with MAE, we follow the same metric called counterfactual effectiveness in Ribeiro et al. (2023) developed by Monteiro et al. (2023), First, trained on the dataset, parent predictors given a value of $x$ can predict parent values, i.e., $(t,i)$’s, and then measure the absolute error between parent values after hard intervention or feasible intervention and their predicted values, which is measured on image the Learned SCM generates given the input of $(t,i)$.

### 20 Feb 2024 v2

@@ 2 Related Work
Counterfactual Explanations. A prominent approach in explainable AI is counterfactual explanation (Wachter et al., 2018; Dhurandhar et al., 2018; Mothilal et al., 2020; Barocas et al., 2020; Verma et al., 2020; Schut et al., 2021), on which our work is likely to have interesting bearings. Most works on this topic define some sense of minimal changes of an input sample with a predicted class $c$ such that adding the minimal changes into the input would make it be classified into another (more desirable) class. Although this paper does not discuss counterfactual explanations, our framework may well be used to define a novel notion of counterfactual explanation by requiring the counterfactual instances to be “natural” in our sense.
@@ 4.3 Distance Measure for Necessary Backtracking
4.4 Identifiability of Natural Counterfactuals
@@ 6.2 MorphoMNIST
MorphoMNIST (Castro et al., 2019) involves three variables $(t,i,x)$. As depicted in Fig. 3 (a), $t$ (digit stroke thickness) causes both $i$ (stroke intensity) and $x$ (images), with $i$ being the direct cause of $x$.
@@
In our experiments, mirroring those in Section 6.1, we incorporate two key changes. First, we utilize two advanced deep learning models, V-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023), for doing counterfactuals. Second, due to the absence of ground-truth SCM for assessing outcome error, we adopt the counterfactual effectiveness metric from (Ribeiro et al., 2023), as developed in (Monteiro et al., 2023). This involves training a predictor on the dataset to estimate parent values $(\hat{t},\hat{i})$ from a counterfactual image $x$ generated by learned model $p(x|t,i)$ with the input $(t,i)$, and then computing the absolute error $|t-\hat{t}|$ or $|i-\hat{i}|$ between the input $(t,i)$ and their predicted counterparts $(\hat{t},\hat{i})$ as inferred from counterfactual images.
@@ 7 Conclusion
To address the impractical scenarios often generated by non-backtracking counterfactuals, we introduces natural counterfactuals. This concept emphasizes real-world applicability by determining feasible intervention, thereby aligning counterfactuals with the data distribution of the actual world. By integrating a naturalness constraint with necessary backtracking, we have developed an optimization framework designed to determine feasible interventions. This method has been effectively demonstrated through various deep learning case studies, highlighting its capability to yield more dependable and precise counterfactuals. Future work involving natural counterfactuals could focus on exploring their applicability and impact in crucial areas like explainable AI and healthcare.
@@ A.2 MorphoMNIST
The MorphoMNIST comes from (Pawlowski et al., 2020), where there are 60000 images as training set and 10,000 images as test dataset. Fig. 8 (a) shows the causal graph for generating Morpho-MNIST; specifically, stroke thickness $t$ causes the brightness intensity $i$, and both thickness $t$ and intensity $i$ cause the digit $x$. Fig. 8 (b) show some samples from Morpho-MNIST. The ground-truth SCMs to generate the Morpho-MNIST are as follows:
@@ Appendix C Model Training
Our study focuses on counterfactual inference and we directly use two state-of-the-art deep-learning SCM models to learn SCMs among variables using a dataset, i.e., D-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023). Specifically, we use code of (Ribeiro et al., 2023) containing the implementation of D-SCM and H-SCM. Take Morpho-MNIST as an example, in both two models, normalizing flows are firstly trained to learn causal mechanisms for all variables except image $x$, i.e., $(t,i)$, and a conditional VAE is used to model image $x$ given its parents $(t,i)$. For D-SCM, the conditional VAE uses normal VAE framework, while H-SCM uses hierarchical VAE structure (Maaløe et al., 2019) to better capture the distribution of images.
@@ H.2.1 MorphoMNIST
We follow the experimental settings of simulation experiments in Sec. 6.1, except for two differences. One is that we use two state-of-the-art deep learning models, namely V-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023), as the backbones to learn counterfactuals. They use normalizing flows to learn causal relationships among $x$’s parent nodes, e.g., $(t,i)$ in MorphoMNIST. Further, to learn $p(x|t,i)$, notice that V-SCM uses VAE (Kingma & Welling, 2014) and HVAE (Maaløe et al., 2019). Another difference is that, instead of estimating the outcome with MAE, we follow the same metric called counterfactual effectiveness in Ribeiro et al. (2023) developed by Monteiro et al. (2023), First, trained on the dataset, parent predictors given a value of $x$ can predict parent values, i.e., $(t,i)$’s, and then measure the absolute error between parent values after hard intervention or feasible intervention and their predicted values, which is measured on image the Learned SCM generates given the input of $(t,i)$.